solve
^3√5x+4+9=4
I just do not understand how to do these kinds of problems.
The 3 is a small square root on the outside

Answer by stanbon(57250)

solve
^3√5x+4+9=4
---
That is not a "small square root"; it is a cube root.
---
Equation:
cbrt(5x+4) + 9 = 4
---
cbrt(5x+4) = -5
---
Raise both sides to the 3rd power to get:
5x+4 = -125
---
5x = -129
---
x = -129/5 = -25 4/5
===========================
Cheers,
Stan H.